# BiMarker: Enhancing Text Watermark Detection for Large Language Models with Bipolar Watermarks

### v1

@@ Abstract
The rapid proliferation of Large Language Models (LLMs) has raised concerns about misuse and the challenges of distinguishing AI-generated text from human-written content. Existing watermarking techniques, such as KGW, still face limitations under low watermark strength, stringent false-positive requirements, and low-entropy scenarios. Our analysis reveals that current detection methods rely on coarse estimates of non-watermarked text, which constrains watermark detectability. We propose the Bipolar Watermark (BiMarker), a novel approach that divides generated text into positive and negative poles, leveraging the difference in green token counts for detection. This differential mechanism significantly enhances the detectability of watermarked text. Theoretical analysis and experimental results demonstrate BiMarker’s effectiveness and compatibility with existing optimization techniques, providing a new dimension for watermarking in LLM-generated content.
Machine Learning, ICML
@@ 1 Introduction
Watermarking techniques offer a solution by enabling LLMs to embed unique, imperceptible identifiers (watermarks) within generated content, distinguishing it from human-written text. A prominent method, KGW [^1] Details can be found in Appendix A.2, achieves high detection performance with low false positive and false negative rates (Kirchenbauer et al., 2023a). KGW works by partitioning the model’s vocabulary into two categories: green and red tokens, with proportions $\gamma$ and $1-\gamma$, respectively. The model is then biased towards green tokens by adjusting their logits with a positive constant (watermark strength), ensuring that the generated text predominantly contains green tokens. During detection, KGW assumes that the number of green tokens in non-watermarked text follows a Gaussian distribution $\mathcal{N}(\mu,\sigma^{2})$, where $\mu=T\cdot\gamma$ and $\sigma^{2}=T\cdot\gamma\cdot(1-\gamma)$, with $T$ representing the number of tokens analyzed. The likelihood that the text is watermarked increases as the observed number of green tokens deviates from the expected $T\cdot\gamma$. In other words, the greater the number of green tokens in the analyzed text exceeds the expected value of $T\cdot\gamma$ for non-watermarked text, the higher the probability that the text is watermarked.
@@ 2.2 Watermarking for LLMs
Watermark embedding during LLM generation can be achieved by modifying token sampling (Kuditipudi et al., 2023; Christ et al., 2023b) or model logits (Kirchenbauer et al., 2023a; Yoo et al., 2024; Ren et al., 2024; Lu et al., 2024), resulting in more natural text. KGW (Kirchenbauer et al., 2023a) represents a promising approach for distinguishing language model outputs from human text while maintaining robustness against realistic attacks, as it can reinforce the watermark with every token. Several studies have built upon KGW, addressing various aspects such as low-entropy generation tasks like code generation (Lee et al., 2024; Lu et al., 2024), enhancing watermark undetectability (Christ et al., 2023a), enhancing performance in payload (Yoo et al., 2024; Wang et al., 2023) and improving robustness (Munyer & Zhong, 2023; Zhao et al., 2023; Liu et al., 2024; Ren et al., 2024). For example,  Fernandez et al. (2023) expanded from a binary red-green partition to a multi-color partition through fine-grained vocabulary partitioning, allowing for the embedding of multi-bit watermarks. Additionally,  Ren et al. (2024) converted semantic embeddings into semantic values using weighted embedding pooling, followed by discretization with NE-Ring, and classified the vocabulary into red-list and green-list categories based on these semantic values, thus enhancing robustness.
@@ 3.1 Movitation
Before presenting our method, we first outline a key assumption: the process of human-written text generation and machine-generated text generation are similar. Specifically, we assume that the process by which a LLM generates text without watermarking can be considered analogous to the process of human text writing. This assumption is grounded in the design of LLMs, which are trained on extensive datasets of human-written text to learn patterns, structures, and styles. As a result, their outputs often demonstrate fluency, coherence, and stylistic consistency that closely resemble human-written text, making this analogy plausible. Thus, by substituting human-written text into (7) with $\alpha=1$, we obtain: $\mathbb{E}|s|_{G}\geq\frac{\gamma T}{1+(\alpha-1)\gamma}S^{\star}=\gamma TS^{\star}$, Where $S^{\star}$ represents the theoretical lower bound. This indicates that the number of green tokens in human-written text is also related to entropy, with the average number of green tokens tending to be lower when entropy is low. Therefore, we believe that during the detection of watermarked text, a lower average number of green tokens suggests a corresponding decrease in the average number of green tokens in human-written text.
We conducted a simple empirical study, assuming that 200 tokens are generated using the KGW watermarking algorithm($\gamma=0.5$ and $\delta=1$, employing polynomial sampling) with OPT-1.3B (Zhang et al., 2022) and prompts sourced from the RealNewsLike subset of the C4 dataset (Raffel et al., 2020) across approximately 500 generations. We detect the number of green tokens in both human-written and watermarked texts. We select the portion of human-written text where the number of green tokens is less than 120, dividing the data into intervals of 2.5 based on the number of green tokens. Then, we calculate the average number of green tokens in human-written text and the corresponding watermarked text for each interval. The relationship between the number of green tokens in watermarked texts and those in human-written texts is analyzed, with the results visualized in Figure 2. From the figure, we observe that when the number of green tokens in human-written text is low, the number of green tokens in watermarked text also tends to be smaller. We calculate their Pearson correlation coefficient, which is 0.700. This demonstrates a strong positive relationship, suggesting that lower green token counts in human-written text are more likely to correspond to lower counts in watermarked text.
@@
However, KGW only consider the theoretical lower bound of green tokens in watermarked texts under varying entropy conditions, overlooking the fact that the expected number of green tokens in non-watermarked texts also varies with entropy. During detection, these methods continue to rely on the difference between the number of green tokens in the analyzed text and the expected fixed value; a larger difference indicates stronger statistical significance. However, this approach is clearly insufficient.
To correct the existing algorithms’ inaccurate estimation of the number of green tokens in human-written texts and its impact on detection results. one intuitive solution would be to utilize a non-watermarked LLM during detection to approximate the number of green tokens in the output text. However, we did not pursue this approach because it would introduce additional inference costs, and in many cases, the prompt itself may not be available. Our goal is to develop an algorithm that does not increase time complexity and can be applied in any scenario.
@@ 3.2 Our Method
From Formula 9, we observe that our numerator still relies on the estimated expectation for the number of green tokens in non-watermarked text. However, this influence is mitigated due to the subtraction of the estimated values for positive and negative polarities. When the number of text sequences generated for the positive and negative polarities satisfies the ratio $T_{p}/T_{n}=(1-\gamma)/\gamma$, the $z$-value simplifies to:
@@ 3.3 Effect of Bipolar Watermark
Consider watermarked text sequences $s$ of $T$ tokens. Each sequence is produced by sequentially sampling a raw probability vector $p(t)$ from the language model, sampling a random green list of size $\gamma|V|$. The sequence $s$ is composed of a positive part $s_{p}$ and a negative part $s_{n}$. The logits for the green and red lists are boosted by $\delta$ before sampling each token in the positive and negative parts, respectively. Define $\alpha=\exp(\delta)$, and let $|s|_{d}$ denote the difference in the number of green list tokens in the positive and negative parts of sequence $s$. We assume that positive and negative tokens are mutually independent, then BiMarker achieves a theoretical lower bound for the $z$-score that consistently exceeds that of KGW.
@@
For non-watermarked text, we assume that the total number of green tokens follows a Gaussian distribution, $\mathcal{N}(\mu_{T},\sigma_{T}^{2})$. Similarly, the numbers of tokens in the positive and negative polarities are assumed to follow Gaussian distributions, $\mathcal{N}(\mu_{T_{p}},\sigma_{T_{p}}^{2})$ and $\mathcal{N}(\mu_{T_{n}},\sigma_{T_{n}}^{2})$, respectively. Moreover, these two distributions are considered independent of each other.
We denote the false positive rate generated by the unipolar detection method as $F_{KGW}$ and the false positive rate generated by the differential detection method as $F_{DIFF}$. Then, we can express the relationship as follows:
@@
SWEET and EWD are advanced algorithms derived from KGW, designed to enhance the statistical significance of watermark text by leveraging entropy. Our proposed algorithm is orthogonal to both of these approaches. The application of our bipolar watermarking and differential detection algorithms, based on SWEET and EWD, is detailed in Appendix B. Moreover, the conclusions of Theorem 3.1 and Theorem 3.2 remain valid when our algorithms are applied to SWEET and EWD. In other words, our algorithm enhances the statistical significance of watermark text in both SWEET and EWD while preserving the original false positive rate. Below, we elaborate on the effectiveness of applying the bipolar watermarking and differential detection algorithms to SWEET, while the rationale for their application to EWD is explained in Appendix B.
@@ 4 Experiments
To assess the performance of the watermark, we primarily analyze the True Positive Rate (TPR) and False Positive Rate (FPR), where TPR represents the detection of watermarked text, and FPR indicates human text falsely flagged as watermarked. We also present the F1 score as an additional metric. Our implementation is based on the completion of KGW with the probability ratio of positive and negative polarities set to $(1-\gamma)/\gamma$.
@@
Our experiments are designed with two scenarios. The first scenario is a high-entropy environment, where we use the same experimental dataset as KGW and employ the OPT-1.3B (Zhang et al., 2022) model to generate watermarked texts. To simulate diverse and realistic language modeling conditions, we randomly select texts from the news-like subset of the C4 dataset (Raffel et al., 2020). Following KGW’s setup, for each randomly selected string, a fixed number of tokens are trimmed from the end to serve as the baseline completion, with the remaining tokens acting as the prompt. In all experiments, we generate 500 samples, each with a length of approximately 200 ± 5 tokens.
@@
Figure 4: Z-scores of Watermarked and Human Texts under Different Parameter Settings. We set the sampling method to multinomial sampling.
@@
Figure 5: The Average Z-Score as a Function of Token Length ($T$) of the Generated Text. Different combinations of hyperparameters were used, and the sampling method was set to multinomial sampling.
@@ 5 Conclusion
Currently, when applying BiMarker with SWEET and EWD in low-entropy scenarios, we adopt a hard-coded approach for polarity assignment. This approach is motivated by the underlying assumption in our bipolarity detection mechanism that positive and negative tokens are independently distributed. However, in the domain of code, syntax and related constraints often result in significant dependencies between nearby tokens, necessitating the use of a hard-coded strategy. Nonetheless, this approach may not ensure watermark integrity in all cases. Additionally, the variation in token counts across different code tasks poses challenges to achieving the theoretically optimal ratio of positive and negative tokens. In future work, we aim to explore more robust methods for polarity assignment to address these limitations.
@@ B.2 Applying BiMarker to EWD
When using bipolar watermarking, we assume that the processes for positive and negative polarities are independent. The number of green tokens in the positive polarity, according to (7), is at least: $\frac{\gamma\alpha T_{p}}{1+(\alpha-1)\gamma}S^{\star}$, Similarly, we assume that the entropy of the negative polarity is maximized, meaning that each token follows a uniform probability distribution. Under this condition, the expected number of green tokens in the negative polarity reaches its maximum, given by: $\frac{\gamma T_{n}}{1+(\alpha-1)\gamma}$, the expected difference between the two polarities of green tokens is given by:
@@
Figure 8: ROC Curves with AUC Values and PPL Under Different Hyperparameter Settings (Normal and Semilog Axes).
@@ D.1 Hyper-parameters
D.2 ROC Curves
Figure 8 presents the ROC curves for KGW and BiMarker with $\gamma$ set to 0.25 and 0.5, and $\delta$ set to $[0.5,0.75,1.0,1.5,2,2.5]$, using polynomial sampling and beam search with a beam size of 8. The perplexity (PPL) of the generated completions is computed using a larger oracle language model (OPT-2.7B). Both standard coordinate axes and logarithmic axes are plotted. The results demonstrate that BiMarker is more effective at distinguishing between watermarked and non-watermarked text, particularly at low false positive rates and low watermark strengths.

### v6

@@ Abstract
The rapid growth of Large Language Models (LLMs) raises concerns about distinguishing AI-generated text from human content. Existing watermarking techniques, like KGW, struggle with low watermark strength and stringent false-positive requirements. Our analysis reveals that current methods rely on coarse estimates of non-watermarked text, limiting watermark detectability. To address this, we propose Bipolar Watermark (BiMarker), which splits generated text into positive and negative poles, enhancing detection without requiring additional computational resources or knowledge of the prompt. Theoretical analysis and experimental results demonstrate BiMarker’s effectiveness and compatibility with existing optimization techniques, providing a new optimization dimension for watermarking in LLM-generated content.
LLMs,Watermarking,Bipolar Watermark,Detectability
@@ 1 Introduction
Figure 1: Illustration of the key idea of BiMarker: Adjust token bias by polarity during generation and detect by the green token difference between polarities.
Watermarking techniques offer a solution by enabling LLMs to embed unique, imperceptible identifiers (watermarks) within generated content, distinguishing it from human-written text (Liu et al., 2024b; Pan et al., 2024b). A prominent method, KGW [^1] Details can be found in Appendix A.2, achieves high detection performance with low false positive and false negative rates (Kirchenbauer et al., 2023a). KGW works by partitioning the model’s vocabulary into two categories: green and red tokens, with proportions $\gamma$ and $1-\gamma$, respectively. The model is then biased towards green tokens by adjusting their logits with a positive constant (watermark strength), ensuring that the generated text predominantly contains green tokens. During detection, KGW assumes that the number of green tokens in non-watermarked text follows a Gaussian distribution $\mathcal{N}(\mu,\sigma^{2})$, where $\mu=T\cdot\gamma$ and $\sigma^{2}=T\cdot\gamma\cdot(1-\gamma)$, with $T$ representing the number of tokens analyzed. The likelihood that the text is watermarked increases as the observed number of green tokens deviates from the expected $T\cdot\gamma$. In other words, the greater the number of green tokens in the analyzed text exceeds the expected value of $T\cdot\gamma$ for non-watermarked text, the higher the probability that the text is watermarked.
@@ 2.2 Watermarking for LLMs
Watermark embedding during LLM generation can be achieved by modifying token sampling (Kuditipudi et al., 2023; Christ et al., 2023b) or model logits (Kirchenbauer et al., 2023a; Yoo et al., 2024; Ren et al., 2024; Lu et al., 2024), resulting in more natural text (Pan et al., 2024a). KGW (Kirchenbauer et al., 2023a) represents a promising approach for distinguishing language model outputs from human text while maintaining robustness against realistic attacks, as it can reinforce the watermark with every token. Several studies have built upon KGW, addressing various aspects such as low-entropy generation tasks like code generation (Lee et al., 2024; Lu et al., 2024), enhancing watermark undetectability (Christ et al., 2023a), enhancing performance in payload (Yoo et al., 2024; Wang et al., 2023) and improving robustness (Munyer & Zhong, 2023; Zhao et al., 2023; Liu et al., 2024a; Ren et al., 2024). For example,  Fernandez et al. (2023) expanded from a binary red-green partition to a multi-color partition through fine-grained vocabulary partitioning, allowing for the embedding of multi-bit watermarks. Additionally,  Ren et al. (2024) converted semantic embeddings into semantic values using weighted embedding pooling, followed by discretization with NE-Ring, and classified the vocabulary into red-list and green-list categories based on these semantic values, thus enhancing robustness.
@@ 3.1 Movitation
Before introducing our method, we first discuss the distribution of human-written text. It is undeniable that the distribution of green tokens in human-written text should be $\gamma T$, but in more fine-grained scenarios, this may not always be accurate. Furthermore, we present our assumption: the expected number of green tokens in human-written text and watermarked text follow a similar trend. In some scenarios, if the expected number of green tokens in the watermarked text is lower, the corresponding human-written text is likely to exhibit a similarly lower expected number of green tokens.
To validate that the number of green tokens in human-written and watermarked texts follows a similar pattern, we conducted a simple empirical study. We used the KGW watermarking algorithm ($\gamma=0.5$, $\delta=1$, and polynomial sampling) with OPT-1.3B (Zhang et al., 2022) to generate 500 watermarked text samples, each consisting of 200 tokens. The prompt content was taken from the RealNewsLike subset of the C4 dataset (Raffel et al., 2020). For each sample from this dataset, we recorded the last 200 tokens, which represent the human-written text under the same prompt as the watermarked text. The remaining tokens served as the prompt.
We selected a subset of the experimental results where the number of green tokens in human-written texts was fewer than 120, grouping them in intervals of 2.5 based on this count. For each group, we computed the average number of green tokens in both the human-written and watermarked texts. The results, shown in Figure 2, demonstrate that when the number of green tokens in human-written text is low, the watermarked text exhibits a similar trend, with fewer green tokens. A Pearson correlation coefficient of 0.7 indicates a strong positive correlation between the two distributions.
@@
However, KGW considers only the theoretical lower bound of green tokens in watermarked text under different entropy conditions, overlooking the fact that the expected number of green tokens in human-written text also varies. During detection, it evaluates the difference between the number of green tokens in the analyzed text and the fixed expected value $\gamma T$, with larger deviations indicating stronger statistical significance. However, treating the number of green tokens in human-written text as a fixed reference value $\gamma T$ is overly simplistic and imprecise.
To overcome the impact of inaccurate estimates of green token distributions in human-written text on detection, we introduce BiMarker. Since prompt content and corresponding human-written text are often unavailable during detection, directly determining the expected number of green tokens for a given prompt is not feasible. Instead, we divide the text under detection into two components—the positive and negative poles (i.e., each token in the text belongs to either the positive or negative pole)—and detect the watermark based on their differences. For instance, when $\gamma=0.5$, we divide the text into two parts with an equal number of tokens. Under ideal partitioning, the expected difference in the number of green tokens between the positive and negative poles in human-written text is zero. This approach incurs no additional inference cost, does not depend on prompt content, and maintains the same time complexity, making it applicable across various scenarios. In the following sections, we detail our method and explain its effectiveness.
@@ 3.2 Our Method
From Formula 1, we observe that our numerator still relies on the estimated expectation for the number of green tokens in non-watermarked text. However, this influence is mitigated due to the subtraction of the estimated values for positive and negative polarities. When the number of text sequences generated for the positive and negative polarities satisfies the ratio $T_{p}/T_{n}=(1-\gamma)/\gamma$, the $z$-value simplifies to:
@@ 3.3 Effect of Bipolar Watermark
Consider watermarked text sequences $s$ of $T$ tokens. Each sequence is produced by sequentially sampling a raw probability vector $p(t)$ from the language model, sampling a random green list of size $\gamma|V|$. The sequence $s$ is composed of a positive part $s_{p}$ and a negative part $s_{n}$. The logits for the green and red lists are boosted by $\delta$ before sampling each token in the positive and negative parts, respectively. Define $\alpha=\exp(\delta)$, and let $|s|_{d}$ denote the difference in the number of green list tokens in the positive and negative parts of sequence $s$. We assume that positive and negative tokens are mutually independent and identically distributed, then BiMarker achieves a theoretical lower bound for the $z$-score that consistently exceeds that of KGW.
@@
We assume that the total number of green tokens in non-watermarked text follows a Gaussian distribution $\mathcal{N}(\mu_{T},\sigma_{T}^{2})$, while the counts of positive and negative tokens follow $\mathcal{N}(\mu_{T_{p}},\sigma_{T_{p}}^{2})$ and $\mathcal{N}(\mu_{T_{n}},\sigma_{T_{n}}^{2})$, respectively, with the distributions being independent. Additionally, we have $T_{p}/T_{n}=(1-\gamma)/\gamma$.
We denote the false positive rates of the unipolar and differential detection methods as $F_{KGW}$ and $F_{DIFF}$, respectively, leading to the relationship:
@@
SWEET and EWD are advanced algorithms derived from KGW, designed to enhance the statistical significance of watermark text by leveraging entropy. Our proposed algorithm is orthogonal to both of these approaches. The application of our bipolar watermarking and differential detection algorithms, based on SWEET and EWD, is detailed in Appendix B. Moreover, the conclusions of Theorem 3.1 and Theorem 3.2 remain valid when our algorithms are applied to SWEET and EWD. In other words, our algorithm enhances the statistical significance of watermark text in both SWEET and EWD while preserving the original false positive rate. Below, we elaborate on the effectiveness of applying the bipolar watermarking and differential detection algorithms to SWEET, while the rationale for their application to EWD is explained in Appendix C.
@@ 4 Experiments
To assess the performance of the watermark, we primarily analyze the True Positive Rate (TPR) and False Positive Rate (FPR), where TPR represents the detection of watermarked text, and FPR indicates human text falsely flagged as watermarked. We also present the F1 score as an additional metric. Furthermore, we provide empirical evidence in Appendix D.2 showing that our watermarking method does not degrade the quality of the generated text. For non-code tasks, we follow Kirchenbauer et al. (2023a) and use use PPL as the evaluation metric. For code generation tasks, we adopt pass@1 accuracy to measure functional correctness, following Lee et al. (2024). Our implementation is based on the completion of KGW with the probability ratio of positive and negative polarities set to $(1-\gamma)/\gamma$.
@@
Our experiments are designed with two scenarios. The first scenario is a high-entropy environment, where we use the same experimental dataset as KGW and employ the OPT-1.3B (Zhang et al., 2022) model to generate watermarked texts. To simulate diverse and realistic language modeling conditions, we randomly select texts from the news-like subset of the C4 dataset (Raffel et al., 2020). Following KGW’s setup, for each randomly selected string, a fixed number of tokens are trimmed from the end to serve as the baseline completion, with the remaining tokens acting as the prompt. We then utilize a larger oracle language model (OPT-2.7B) to compute the perplexity (PPL) of the generated completions. In all experiments, we generate 500 samples, each with a length of approximately 200 ± 5 tokens.
@@
Figure 4: Z-scores of Watermarked and Human Texts under Different Parameter Settings with multinomial sampling.
@@
Figure 5: The Average Z-Score as a Function of Token Length ($T$) of the Generated Text with multinomial sampling.
@@ 5 Conclusion
BiMarker uses hard-coded polarity assignment with SWEET and EWD in low-entropy scenarios, assuming independent token distributions. However, code syntax introduces adjacent token dependencies—we mitigate this by enforcing minimum spacing between opposite polarities, though watermark integrity risks persist. Varying token counts also hinder optimal polarity ratios. Future work will refine polarity assignment robustness.
@@ B.2 Applying BiMarker to EWD
When using bipolar watermarking, the number of green tokens in the positive polarity, according to (7), is at least: $\frac{\gamma\alpha T_{p}}{1+(\alpha-1)\gamma}S^{\star}$, Similarly, we assume that the entropy of the negative polarity is maximized, meaning that each token follows a uniform probability distribution. Under this condition, the expected number of green tokens in the negative polarity reaches its maximum, given by: $\frac{\gamma T_{n}}{1+(\alpha-1)\gamma}$, the expected difference between the two polarities of green tokens is given by:
@@ D.1 Hyper-parameters
D.2 Evaluating the Impact of BiMarker on Text Quality
Figure 8 presents a PPL comparison under different watermark strengths, where the x-axis denotes BiMarker’s PPL and the y-axis shows the ratio of KGW’s PPL to BiMarker’s. We observe that, under the same watermark strength, BiMarker and KGW exhibit similar PPL values, indicating that BiMarker does not significantly alter the PPL compared to BiMarker, which aligns with our theoretical prediction.
Table 4 presents the impact of applying BiMarker to various watermark embedding algorithms on the quality of generated code. Note that EWD does not include a generation algorithm. The results clearly indicate that BiMarker leaves code generation quality essentially unchanged.
Figure 8: PPL comparison under different watermark strengths ($\delta\in[0.5,0.75,1.0,1.5,2,2.5]$), where the x-axis denotes BiMarker’s PPL and the y-axis shows the ratio of KGW’s PPL to BiMarker’s.
Table 4: The accuracy (pass@1) of the generated code using different methods($\gamma=0.25,\delta=2$).
methods humaneval mbpp
KGW 0.274 0.326
KGW+BiMarker 0.287 0.336
SWEET 0.311 0.334
SWEET+BiMarker 0.323 0.31